5. In a recent year the magnitudes​ (Richter scale) of​ 10,594 earthquakes were recorded. The mean is 1.268 and the standard deviation is 0.571. Consider the magnitudes that are unusual. What are the magnitudes that separate the unusual magnitudes from those that are​ usual? ​(Consider a value to be unusual if its z score is less than minus2 or greater than​ 2.)

The lowest bounds earthquakes magnitude is ___

The highest bounds earthquakes magnitude is ___

9.Below are 36 sorted ages of an acting award winner. Find the percentile corresponding to age 70 using the method presented in the textbook.

16
17
17
20
21
22
24
26
26
30
32
37
39
44
45
46
48
50
51
57
57
60
60
65
65
66
68
68
70
71
73
73
74
77
78
79

percentile of value 70=?
Answer #1

Solution:-

5)

The lowest bounds earthquakes magnitude is 1.257.

The highest bounds earthquakes magnitude is 1.279.

Mean = 1.268, S.D = 0.571

C.I = \bar{x}\pm 2\ast \frac{\sigma }{\sqrt{n}}

C.I = 1.268 + 2*0.0055476

C.I = 1.268 + 0.011095

C.I = (1.257, 1.279)